Find the real interest rate (when the compound interest is quarterly) when a capital of EUR 29,000 gives a final value of EUR 33,000 in 6 years. WITHOUT Excel please.
Answer #1
Solution:
Real interest rate is
Real rate of interest = 2.15934% in 5 decimal
Real rate of interest = 2.16% in 2 decimal
Use as per requirement of question , but remember when you take 5 or 6 decimal then you will get exact results not in two decimal as the question is decimal sensitive.
Working Notes:
Real rate of interest is the rate at which EUR 29000 is compounded quarterly to future value in 6 years to EUR 33,000 , Hence we have to used concept future value of present value of deposit to get the annual rate.
C0 = Capital= Present value = $29000
FV=Future value of the deposit after 6 years =$33000
r= rate per period=Quarterly interest rate = r=(Real rate of interest /4)=??
n = no. Of periods = No. of quarter in a year x no. of years = 4 x 6 =24
Using future value formula
FV= C0 x (1+r)^n
33000 = 29000 (1+ r)^24
(33000/29000) = (1+r)^24
Taking Log on both side
Log(33000/29000) = Log(1+r)^24
Log(1.13793103448) = Log(1+r)^24
using relation loga^b = b x Log a
Log(1.13793103448) = 24 x Log(1+r) Log(1.13793103448) = 0.056115942
(0.056115942) =24 x Log (1+r)
(0.056115942/24) = Log (1+r)
(0.002338164249) = Log (1+r)
now we move log from right side to left side then it becomes antilog
antilog(0.002338164249) = (1+r)
(1+r) = antilog (0.002338164249)
(1+r) = antilog (0.002338164249) We got from online financial calculator
(1+r) = 1.0053983409590392 antilog(0.002338164249) =1.0053983409590392
r= 1.0053983409590392- 1
r= 0.0053983409590392
r= 0.53983409590%
r= Real rate of interest/4 =0.53983409590%
Real rate of interest = 0.53983409590% x 4
Real rate of interest = 2.1593363836%
Real rate of interest = 2.15934% in 5 decimal
Real rate of interest = 2.16% in 2 decimal
Lets Check is our computed rate is correct.
Using future value formula
FV= C0 x (1+r)^n
33000 = 29000 x (1+ r)^24
33000 = 29000 x (1+(2.15934%/4))^24
33000 = 33000.00712
33000 =33000
Hence above computed rate is correct only you uses 5 decimal or above.
Please feel free to ask if anything about above solution in comment section of the question.